# SAIF: Sparse Adversarial and Imperceptible Attack Framework

### v3

@@ Abstract
Adversarial attacks hamper the decision-making ability of neural networks by perturbing the input signal. For instance, adding calculated small distortions to images can deceive a well-trained image classification network. In this work, we propose a novel attack technique called Sparse Adversarial and Imperceptible Attack Framework (SAIF). Specifically, we design imperceptible attacks that contain low-magnitude perturbations at a few pixels and leverage these sparse attacks to reveal the vulnerability of classifiers. We use the Frank-Wolfe (conditional gradient) algorithm to simultaneously optimize the attack perturbations for bounded magnitude and sparsity with $O(1/\sqrt{T})$ convergence. Empirical results show that SAIF computes highly imperceptible and interpretable adversarial examples, and largely outperforms state-of-the-art sparse attack methods on ImageNet and CIFAR-10.
@@ 1 Introduction
  • We show through comprehensive experiments that, for tight sparsity and magnitude constraints, SAIF outperforms state-of-the-art sparse attacks by a large margin (by 2$\geq$ 2$\times$ 2 × higher fooling rates for most thresholds).
@@ 2 Related Works
The Jacobian-based Saliency Map Attack (JSMA) denotes pixel-wise saliency by backpropagated gradient magnitudes, then searches over the most salient pixels for a sparse targeted perturbation (Papernot et al., 2016). This attack is slow and fails to scale to larger images. SparseFool (Modas et al., 2019) extends DeepFool (Moosavi-Dezfooli et al., 2016) to a sparse attack within the valid pixel magnitude bounds. The attack, however, is untargeted. Croce & Hein (2019) devise a black-box attack by evaluating the impact of each pixel on logits and randomly sampling salient pixels to find a feasible sparse combination. Similar to JSMA, these attacks are expensive to compute, visually noticeable, and unstructured. The same limitations hold for SA-MOO (Williams & Li, 2023). StrAttack (Xu et al., 2019) uses ADMM (Alternating Direction Method of Multipliers) to optimize for group sparsity and perturbation magnitude constraints, but the perturbations are computationally expensive, visually noticeable, and of low sparsity. SAPF (Fan et al., 2020) also uses ADMM with projections to solve a factorized objective. It fails to converge for a tighter sparsity budget, requires extensive hyperparameter tuning, and has a prohibitively long running time. Among generator-based methods, Dong et al. (2020) propose GreedyFool, a two-stage approach to greedily sparsify perturbations obtained from a generator. Similarly, TSAA (He et al., 2022) generates sparse, magnitude-constrained adversarial attacks with high black-box transferability. The perturbations are typically spatially contiguous and are therefore more noticeable than other sparse attacks. The design of our attack, and employing Frank-Wolfe for optimization, yield highly sparse and inconspicuous adversarial examples efficiently.
@@
A fairly novel research direction examines adversarial examples and model explanations in conjunction, by noting the overlap between core ideas in the domains. On simple datasets such as MNIST, Ignatiev et al. (2019) demonstrates a hitting set duality between model explanations and adversarial examples. Similarly, Wang et al. (2022) leverage adversarial attacks to devise a novel model explainer. Xu et al. (2019) examine the correspondence of attack perturbations with discriminative image features. We formulate our attack with an explicit sparsity constraint which emphasizes only the most vulnerable pixels in an image. We also empirically analyze the overlap of adversarial perturbations and salient regions in images.
@@ 3.1 Adversarial Attacks
Given an image $\mathbf{x}\in\mathbb{R}^{h\times w\times c}$, a trained classifier $f:\mathbb{R}^{h\times w\times c}\xrightarrow{}\{1\ldots k\}$ that maps the image to one of $k$ classes, and $f(\mathbf{x})=c$. Adversarial attacks aim at finding $\mathbf{x}^{\prime}$ that is very similar to $\mathbf{x}$ by a distance metric, i.e. $||\mathbf{x}-\mathbf{x}^{\prime}||_{p}\leq\epsilon$, $(p\in\{0,1,2,\infty\},$ $\epsilon$ is small) such that $f(\mathbf{x}^{\prime})=t$, where $t\neq c$.
Depending on the adversary’s knowledge of the target model, adversarial attacks can be white-box (known model architecture and parameters) or black-box (unknown learning algorithm, the attacker only sees the most likely prediction given an input). We adopt the white-box setting in this work, however, gradient approximations (Chen et al., 2020) can be used for black-box attacks.
@@ 4 Method
$\max_{\mathbf{s},\mathbf{p}}D(\mathbf{s,p})_{adv},\ \text{s.t.}\|\mathbf{s}\|_{1}\leq k,\ \mathbf{s}\in[0,1]^{h\times w\times c},\|\mathbf{p}\|_{\infty}\leq\epsilon$ (2)
@@
$\min_{\mathbf{s},\mathbf{p}}D(\mathbf{s,p})_{\tilde{c},adv},\ \text{s.t.}\|\mathbf{s}\|_{1}\leq k,\ \mathbf{s}\in[0,1]^{h\times w\times c},\|\mathbf{p}\|_{\infty}\leq\epsilon$ (4)
@@
To constrain $\mathbf{s}$ we use a non-negative $k$-sparse polytope, which is a convex hull of the set of vectors in $[0,1]^{h\times w\times c}$, each vector admitting at most $k$ non-zero elements. We adopt the method in Macdonald et al. (2022) to perform the LMO over this polytope. That is, for $\mathbf{z_{t}}$ we choose the vector with at most $k$ non-zero entries, where the conditional gradient $\textbf{m}_{t}^{s}$ assumes $k$ smallest negative values (thus highest in magnitude). These $k$ components of $\mathbf{z_{t}}$ are then set to 1 and the rest to zero. For example, if $k=10$ and there are 20 negative values in $\textbf{m}_{t}^{s}$, the 10 smallest values are set to 1 and the rest to 0.
@@
Since the objective of SAIF is non-convex, a monotonicity guarantee is helpful to ensure that the separate optimizations of each variable sync well. To this end, we use the following adaptive step size formulation (Carderera et al., 2021; Macdonald et al., 2022) for monotonicity in the objective:
@@
Input: Clean image $\mathbf{x}\in[0,I_{max}]^{h\times w\times c}$, target class $\tilde{c}\in$ ${1\ldots k}$, $\textbf{s}_{0}\in\mathcal{C}_{s}=\{\textbf{s}\in[0,1]^{h\times w\times c}:\|\textbf{s}\|_{1}\leq k\}$, $\textbf{p}_{0}\in\mathcal{C}_{p}=\{\textbf{p}\in[0,I_{max}]^{h\times w\times c}:\|\textbf{p}\|_{\infty}\leq\epsilon\}$.
@@ 3      $\textbf{m}^{s}_{t}=\nabla_{s}D(\textbf{s}_{t-1},\textbf{p}_{t-1})$
4      $_{t}=\text{argmin}_{\textbf{v}\in\mathcal{C}_{p}}\left<\textbf{m}^{p}_{t},\textbf{v}\right>$ v ⟩
5      $_{t}=\text{argmin}_{\textbf{z}\in\mathcal{C}_{s}}\left<\textbf{m}^{s}_{t},\textbf{z}\right>$ z ⟩
@@ 5 Experiments
We use the ImageNet classification dataset (ILSVRC2012) (Krizhevsky et al., 2012) in our experiments, which has $[299\times 299]$ RGB images belonging to 1,000 classes. We evaluate all attacks on 5,000 samples chosen from the validation set. For classification, we test on two deep convolutional neural network architectures, namely Inception-v3 (top-1 accuracy: 77.9%) and ResNet-50 (top-1 accuracy: 74.9%). We use the pre-trained models from Keras applications (Chollet et al., 2015). We also report results on CIFAR-10 in the appendix.
@@ 5.1 Evaluation Metrics
  • Attack Success Rate (ASR). A targeted adversarial attack is deemed successful if perturbing the image fools the classifier into labeling it with a premeditated target class $\tilde{c}$. An untargeted attack is successful if it leads the classifier to predict any incorrect class. Given $n$ images in a dataset, if $m$ attacks are successful, the attack success rate is defined as ASR $=m/n(\%)$.
Note that for RGB images, SparseFool (Modas et al., 2019) and Greedyfool (Dong et al., 2020) average the perturbation $\mathbf{p}$ across the channels and report the ASR for sparsity $||\mathbf{p}_{\text{flat}}||_{0}\leq k$, where $\mathbf{p}_{\text{flat}}\in[0,I_{max}]^{h\times w}$. Since for SAIF the sparsity constraint $k$ applies across all $h\times w\times c$ pixels, we report the ASR for all methods without averaging the final perturbation across the channels (i.e. for $||\mathbf{p}||_{0}\leq k$, $\mathbf{p}\in[0,I_{max}]^{h\times w\times c}$).
@@
Table 1: Quantitative evaluation of targeted attack on ImageNet. We report the ASR for varying constraints on sparsity ‘$k$’ and $\ell_{\infty}$-norm of the magnitude of perturbation ‘$\epsilon$’. (* TSAA codebase lacks pre-trained generators for targeted attacks on ResNet50 and lower $\epsilon$.)
@@
    In the event of perfect correspondence between the GT salient regions and adversarial perturbation, Loc.$\rightarrow$1. Whereas, Loc.$\rightarrow$0 when there is poor overlap between the two.
@@ 6 Results
Table 2: Quantitative evaluation of untargeted attack on Inception-v3 and ResNet50 trained on ImageNet dataset. We report the ASR for varying constraints on sparsity ‘$k$’ and $\ell_{\infty}$-norm of the magnitude of perturbation ‘$\epsilon$’. Black-box attacks are marked with *.
@@
Table 1 demonstrates that SAIF consistently outperforms both targeted attack baselines by a large margin for all $\epsilon$ and $k$ thresholds. For smaller $\epsilon$, GreedyFool fails to attack samples unless the sparsity threshold is significantly relaxed. A similar pattern is observed for the sparsity budget - competing attacks completely fail for tighter bounds on perturbation magnitude (see $\epsilon=10/100/255$ at $k=2000$). Moreover, at lower $k$, ResNet50 is easier to fool than Inception-v3.
We also compare the ASR for untargeted attacks against the baselines in Table 2. Here as well SAIF significantly outperforms other attacks, particularly on tighter perturbation bounds. By jointly optimizing over the two constraints, we are able to fool DNNs with extremely small distortions. For instance, at $\epsilon=10$, SAIF modifies only 0.37% pixels per image to successfully perturb 89.02% of input samples. This is more than twice the ASR of state-of-the-art sparse attack methods. Similarly, at $\epsilon=255$, only 0.03% pixels are attacked to achieve a perfect ASR. On CIFAR-10, SAIF achieves $\sim$3$\times$ higher ASR on lower sparsity thresholds. The results are reported in the Appendix.
@@
Note that sparse attacks like ours address a more challenging problem than $\ell_{\infty}$-norm threat models since both the perturbation magnitude and sparsity are to be constrained. For $\ell_{\infty}$ constraints ϵ=italic-$\epsilon=$10, 100, and 255, AutoAttack (Croce & Hein, 2020) achieves $100\%$ ASR while perturbing $\sim$98of image pixels for each $\epsilon$. For the same $\epsilon$, SAIF achieves $100\%$ ASR with 1.12%, 0.07%, and 0.037% sparsity, respectively. Moreover, sparse attacks are more visually interpretable than $\ell_{\infty}$ attacks.
@@
For untargeted attacks, we present examples from all competing attack methods in Figure 3. We ease the magnitude constraint to allow all baselines to achieve some successful attacks at the same level of sparsity. It can be observed that all competing methods produce highly conspicuous perturbations around the face of the leopard and on the outlines of the forklift and the albatross. PGD $\ell_{\infty}+\ell_{0}$, in particular, adds the most noticeable distortions to images. In contrast, SAIF attack produces adversarial examples that appear virtually identical to the clean input images.
@@ 6.4 Comparison against non-sparse methods
To emphasize the significance of an explicit sparsity constraint for adversarial attacks, we report the percentage of pixels perturbed by three non-sparse attacks against SAIF. Comprehensive results are in table 4. SAIF achieves perfect ASR by modifying $\leq$1of pixels. The same ASR is achieved by non-sparse baselines by modifying $\sim 99\%$ pixels for all constraints on perturbation magnitude.
@@ 6.5 Speed Comparison
Table 5 reports running times of all baselines. PGD $\ell_{0}+\ell_{\infty}$ (Croce & Hein, 2019) runs the fastest but produces the most noticeable perturbations. Note that, although the inference time for GreedyFool (Dong et al., 2020) and TSAA (He et al., 2022) is $\leq 2$ seconds, these generator-based attacks require pre-training a generator for each target model (as well as each $\epsilon$ and target class for TSAA (He et al., 2022)). This incurs a significant computational overhead ($>$7 days on a single GPU), which offsets their faster optimization times. SAIF attack relies solely on pre-trained classifiers, and is significantly faster than the existing state-of-the-art Homotopy-Attack (Zhu et al., 2021). Moreover, more efficient implementations of the FW LMO can further shorten running times, which we leave for future work.

### v4

@@ Abstract
Adversarial attacks hamper the decision-making ability of neural networks by perturbing the input signal. For instance, adding calculated small distortions to images can deceive a well-trained image classification network. In this work, we propose a novel attack technique called Sparse Adversarial and Imperceptible Attack Framework (SAIF). Specifically, we design imperceptible attacks that contain low-magnitude perturbations at a few pixels and leverage these sparse attacks to reveal the vulnerability of classifiers. We use the Frank-Wolfe (conditional gradient) algorithm to simultaneously optimize the attack perturbations for bounded magnitude and sparsity with $O(1/\sqrt{T})$ convergence. Empirical results show that SAIF computes highly imperceptible and interpretable adversarial examples, and largely outperforms state-of-the-art sparse attack methods on ImageNet and CIFAR-10. Implementation of SAIF is available at https://github.com/toobaimt/SAIF.
@@ 1 Introduction
  • We show through comprehensive experiments that, for tight sparsity and magnitude constraints, SAIF outperforms state-of-the-art sparse attacks by a large margin (by $\geq 2\times$ higher fooling rates for most thresholds).
@@ 2 Related Works
The Jacobian-based Saliency Map Attack (JSMA) denotes pixel-wise saliency by backpropagated gradient magnitudes, then searches over the most salient pixels for a sparse targeted perturbation (Papernot et al., 2016). This attack is slow and fails to scale to larger images. SparseFool (Modas et al., 2019) extends DeepFool (Moosavi-Dezfooli et al., 2016) to a sparse attack within the valid pixel magnitude bounds. The attack, however, is untargeted. Croce & Hein (2019) devise a black-box attacks PGD $\ell_{0}+\ell_{\infty}$ and PGD $\ell_{0}+\sigma$ by evaluating the impact of each pixel on logits and randomly sampling salient pixels to find a feasible sparse combination. Similar to JSMA, these attacks are expensive to compute, visually noticeable, and unstructured. The same limitations hold for SA-MOO (Williams & Li, 2023). StrAttack (Xu et al., 2019) uses ADMM (Alternating Direction Method of Multipliers) to optimize for group sparsity and perturbation magnitude constraints, but the perturbations are computationally expensive, visually noticeable, and of low sparsity. SAPF (Fan et al., 2020) also uses ADMM with projections to solve a factorized objective. It fails to converge for a tighter sparsity budget, requires extensive hyperparameter tuning, and has a prohibitively long running time. Among generator-based methods, Dong et al. (2020) propose GreedyFool, a two-stage approach to greedily sparsify perturbations obtained from a generator. Similarly, TSAA (He et al., 2022) generates sparse, magnitude-constrained adversarial attacks with high black-box transferability. The perturbations are typically spatially contiguous and are therefore more noticeable than other sparse attacks. The design of our attack, and employing Frank-Wolfe for optimization, yield highly sparse and inconspicuous adversarial examples efficiently.
@@
A fairly novel research direction examines adversarial examples and model explanations in conjunction, noting the overlap between core ideas in the domains. On simple datasets such as MNIST, Ignatiev et al. (2019) demonstrates a hitting set duality between model explanations and adversarial examples i.e. there exists a tight overlap between features identified as important for a model’s prediction and those targeted by successful adversarial attacks (and that one can be recovered from the other), suggesting that both explanations and attacks often focus on the same critical input features. Similarly, Wang et al. (2022) leverage adversarial attacks to devise a novel model explainer. Xu et al. (2019) examine the correspondence of attack perturbations with discriminative image features. We formulate our attack with an explicit sparsity constraint, which emphasizes only the most vulnerable pixels in an image. We also empirically analyze the overlap of adversarial perturbations and salient regions in images.
@@ 3.1 Adversarial Attacks
Given an image $\mathbf{x}\in\mathbb{R}^{h\times w\times ch}$, a trained classifier $f:\mathbb{R}^{h\times w\times{ch}}\xrightarrow{}\{1\ldots C\}$ that maps the image to one of $C$ classes, and $f(\mathbf{x})=c$. Adversarial attacks aim at finding $\mathbf{x}^{\prime}$ that is very similar to $\mathbf{x}$ by a distance metric, i.e. $\mathbf{x}-\mathbf{x}^{\prime}||_{p}\leq\epsilon$, $(p\in\{0,1,2,\infty\},$ $\epsilon$ is small) such that $f(\mathbf{x}^{\prime})=t$, where $t\neq c$.
Depending on the adversary’s knowledge of the target model, adversarial attacks can be white-box (known model architecture and parameters) or black-box (unknown learning algorithm; the attacker only sees the most likely prediction given an input). Whether SAIF can be extended to the black-box setting (via, e.g., gradient approximations (Chen et al., 2020)) is not addressed in this work and deserves study.
@@ 4 Method
$\max_{\mathbf{s},\mathbf{p}}D(\mathbf{s,p})_{adv},\ \text{s.t.}\|\mathbf{s}\|_{1}\leq k,\ \mathbf{s}\in[0,1]^{h\times w\times ch},\|\mathbf{p}\|_{\infty}\leq\epsilon$ (2)
@@
$\min_{\mathbf{s},\mathbf{p}}D(\mathbf{s,p})_{\tilde{c},adv},\ \text{s.t.}\|\mathbf{s}\|_{1}\leq k,\ \mathbf{s}\in[0,1]^{h\times w\times ch},\|\mathbf{p}\|_{\infty}\leq\epsilon$ (4)
@@
To constrain $\mathbf{s}$ we use a non-negative $k$-sparse polytope, which is a convex hull of the set of vectors in $0,1]^{h\times w\times{\color[rgb]{1,0,0}ch}}$, each vector admitting at most $k$ non-zero elements. We adopt the method in Macdonald et al. (2022) to perform the LMO over this polytope. That is, for $\mathbf{z_{t}}$ we choose the vector with at most $k$ non-zero entries, where the conditional gradient $\textbf{m}_{t}^{s}$ assumes $k$ smallest negative values (thus highest in magnitude). These $k$ components of $\mathbf{z_{t}}$ are then set to 1 and the rest to zero. For example, if $k=10$ and there are 20 negative values in $\textbf{m}_{t}^{s}$, the 10 smallest values are set to 1 and the rest to 0.
@@
Since the objective of SAIF is non-convex, a monotonicity guarantee is helpful to ensure that the separate optimizations of each variable sync well. Such monotonicity guarantee facilitates coordinated convergence of the sparse mask and perturbation variables, ensuring that both sparsity and magnitude constraints are jointly optimized in a stable and synchronized manner. To this end, we use the following adaptive step size formulation (Carderera et al., 2021; Macdonald et al., 2022) to ensure monotonicity in the objective , which, together with the iterates being convex combinations, ensures that solutions do not leave the domain of feasible solutions:
@@
Input: Clean image $\mathbf{x}\in[0,I_{max}]^{h\times w\times c}$, $\textbf{s}_{0}\in\mathcal{C}_{s}=\{\textbf{s}\in[0,1]^{h\times w\times ch}:\|\textbf{s}\|_{1}\leq k\}$, $\textbf{p}_{0}\in\mathcal{C}_{p}=\{\textbf{p}\in[0,I_{max}]^{h\times w\times ch}:\|\textbf{p}\|_{\infty}\leq\epsilon\}$.
@@ 3 $\textbf{m}^{s}_{t}=\nabla_{s}D(\textbf{s}_{t-1},\textbf{p}_{t-1})$
4 $\textbf{v}_{t}=\text{argmin}_{\textbf{v}\in\mathcal{C}_{p}}\left<\textbf{m}^{p}_{t},\textbf{v}\right$>
5 $\textbf{z}_{t}=\text{argmin}_{\textbf{z}\in\mathcal{C}_{s}}\left<\textbf{m}^{s}_{t},\textbf{z}\right$>
@@ 5 Experiments
We use the ImageNet classification dataset (ILSVRC2012) (Krizhevsky et al., 2012) in our experiments, which has $299\times 299]$ RGB images belonging to 1,000 classes. We evaluate all attacks on 5,000 samples chosen from the validation set. For classification, we test on two deep convolutional neural network architectures, namely Inception-v3 (top-1 accuracy: 77.9%) and ResNet-50 (top-1 accuracy: 74.9%). We use the pre-trained models from Keras applications (Chollet et al., 2015). We also report results on CIFAR-10 in the appendix.
@@ 5.1 Evaluation Metrics
  • Attack Success Rate (ASR). A targeted adversarial attack is deemed successful if perturbing the image fools the classifier into labeling it with a premeditated target class $\tilde{c}$. An untargeted attack is successful if it leads the classifier to predict any incorrect class. Given $n$ images in a dataset, if $m$ attacks are successful, the attack success rate is defined as ASR $m/n(\%)$.
Note that for RGB images, SparseFool (Modas et al., 2019) and Greedyfool (Dong et al., 2020) average the perturbation $\mathbf{p}$ across the channels and report the ASR for sparsity $\mathbf{p}_{\text{flat}}||_{0}\leq k$, where $\mathbf{p}_{\text{flat}}\in[0,I_{max}]^{h\times w}$. Since for SAIF the sparsity constraint $k$ applies across all $h\times w\times{\color[rgb]{1,0,0}ch}$ pixels, we report the ASR for all methods without averaging the final perturbation across the channels (i.e. for $\mathbf{p}||_{0}\leq k$, $\mathbf{p}\in[0,I_{max}]^{h\times w\times{\color[rgb]{1,0,0}ch}}$).
@@
Table 1: Quantitative evaluation of targeted attack on ImageNet. We report the ASR for varying constraints on sparsity ‘$k$’ and $\ell_{\infty}$-norm of the magnitude of perturbation $\epsilon$’. (* TSAA codebase lacks pre-trained generators for targeted attacks on ResNet50 and lower $\epsilon$.)
@@ 6 Results
Table 2: Quantitative evaluation of untargeted attack on Inception-v3 and ResNet50 trained on ImageNet dataset. We report the ASR for varying constraints on sparsity ‘$k$’ and $\ell_{\infty}$-norm of the magnitude of perturbation $\epsilon$’. Black-box attacks are marked with *.
@@
Table 1 demonstrates that SAIF consistently outperforms both targeted attack baselines by a large margin for all $\epsilon$ and $k$ thresholds. For smaller $\epsilon$, GreedyFool fails to attack samples unless the sparsity threshold is significantly relaxed. A similar pattern is observed for the sparsity budget - competing attacks completely fail for tighter bounds on perturbation magnitude (see $\epsilon\in\{10,100,255\}$ at $k=2000$). Moreover, at lower $k$, ResNet50 is easier to fool than Inception-v3.
[…]
@@
Note that sparse attacks like ours address a more challenging problem than $\ell_{\infty}$-norm threat models since both the perturbation magnitude and sparsity are to be constrained. For $\ell_{\infty}$ constraints $\epsilon=$10, 100, and 255, AutoAttack (Croce & Hein, 2020) achieves $100\%$ ASR while perturbing $\sim$98% of image pixels for each $\epsilon$. For the same $\epsilon$, SAIF achieves $100\%$ ASR with 1.12%, 0.07%, and 0.037% sparsity, respectively. Moreover, sparse attacks are more visually interpretable than $\ell_{\infty}$ attacks.
We additionally evaluate SAIF against the One Pixel Attack (Su et al., 2019), which aims to fool classifiers by modifying only a single pixel per image. For ResNet50 trained on ImageNet, the One Pixel Attack is reported to achieve 34.0% ASR for untargeted and 14.4% ASR for targeted settings. Despite not being specifically optimized for such an extreme sparsity regime, SAIF achieves a fairly competitive 20.93% ASR in the untargeted setting and 6.04% in the targeted case, demonstrating its robustness even under highly constrained attack scenarios.
@@
For untargeted attacks, we present examples from all competing attack methods in Figure 3. We ease the magnitude constraint to allow all baselines to achieve some successful attacks at the same level of sparsity. Upon a closer visual inspection, it can be observed that competing methods produce significantly more conspicuous perturbations around the face of the leopard and on the outlines of the forklift and the albatross. PGD $\ell_{\infty}+\ell_{0}$, in particular, adds the most noticeable distortions to images. In contrast, SAIF attack produces adversarial examples that appear virtually identical to the clean input images.
@@ 6.4 Comparison against non-sparse methods
To emphasize the significance of an explicit sparsity constraint for adversarial attacks, we report the percentage of pixels perturbed by three non-sparse attacks against SAIF. Comprehensive results are in table˜4. SAIF achieves perfect ASR by modifying $\leq$1% of pixels. The same ASR is achieved by non-sparse baselines by modifying $\sim 99\%$ pixels for all constraints on perturbation magnitude.
@@ 6.5 Speed Comparison
Table 5 reports running times of all baselines. PGD $\ell_{0}+\ell_{\infty}$ (Croce & Hein, 2019) runs the fastest but produces the most noticeable perturbations. Note that, although the inference time for GreedyFool (Dong et al., 2020) and TSAA (He et al., 2022) is $\leq 2$ seconds, these generator-based attacks require pre-training a generator for each target model (as well as each $\epsilon$ and target class for TSAA (He et al., 2022)). This incurs a significant computational overhead (>>7 days on a single GPU), which offsets their faster optimization times. SAIF attack relies solely on pre-trained classifiers, and is significantly faster than the existing state-of-the-art Homotopy-Attack (Zhu et al., 2021). Moreover, more efficient implementations of the FW LMO can further shorten running times, which we leave for future work.